Suppose that you hold a piece of land in the City of London that you may want to sell in one year. As a U.S. resident, you are concerned with the dollar value of the land. Assume that, if the British economy booms in the future, the land will be worth £2,905 and one British pound will be worth $1.42. If the British economy slows down, on the other hand, the land will be worth less, i.e., £1,705, but the pound will be stronger, i.e., $1.64/£. You feel that the British economy will experience a boom with a 60% probability and a slow-down with a 40% probability.

(a) Estimate your exposure b to the exchange risk.
Answer #1

Convert the price into dollars

2905*1.42 = $4125.1

1705*1.64= $2796.2

Expected price

E(P) = 0.6*4125.1 + 0.4*2796.2 = 3,593.54

E(S) = 0.6*1.42+ 0.4*1.64=1.508

Variance = (0.6)(1.508-1.42)2 + (0.4)(1.508-1.64)2 = 0.0116

Covariance, Cov(P,S) = 0.6(4125.1-3593.54)(1.42-1.508) + (0.4)(2796.2-3593.54)(1.64-1.508) = -70.16

b = cov(P,S)/Var(S) = -70.16/0.0116 = -6,048.3R.

Since the exposure is negative, as the pound value increases against dollar, value of the British holding decreases.